Determine the linear function for a graph that is a line containing the points
(-1,-1) and (3,-9)

Answer by stanbon(75887)

Determine the linear function for a graph that is a line containing the points
(-1,-1) and (3,-9)
---------------------
slope = (-9--1)/(3--1) = -8/4 = -2
-----
Equation Form:
y = mx + b
solve for "b":
-1 = -2(-1)+b
-1 = 2 = b
b = -3
-----
Equation:
y = -2x - 3
================
Cheers,
Stan H.